OpenStudy (amenah8):

OPTIMIZATION: According to U.S. postal regulations, the girth plus the length of a parcel sent by mail may not exceed 108 inches, where by "girth" we mean the perimeter of the smallest end. What is the largest possible volume of a rectangular parcel with a square end that can be sent by mail? Such a package is shown below. Assume y>xy>x. What are the dimensions of the package of largest volume? Find a formula for the volume of the parcel in terms of x and y. Volume =?

OpenStudy (agent0smith):

Volume will be something like \(\large V = x^2y\) if x is the dimension of the square base. The length plus girth would be \(\large 108 = 4x+y\), since the perimeter (girth) of the square end is 4x. Solve the second eq. for y, plug it into the first. Differentiate, set equal to zero, solve. You might want to use a first or second derivative test to check it's a maximum. Then plug the x value back into the volume equation, to work out the maximum volume.

OpenStudy (bonnieisflash1.0):

need help @Amenah8

OpenStudy (amenah8):

I'm trying to do it based on what @agent0smith said but I'm still a bit confused

OpenStudy (bonnieisflash1.0):

well almost all of my people are offline

OpenStudy (amenah8):

i just messaged someone

OpenStudy (bonnieisflash1.0):

i would call @zepdrix but he is offline

OpenStudy (bonnieisflash1.0):

@Jamierox4ev3r help

OpenStudy (bonnieisflash1.0):

do you know someone smart in your fan

OpenStudy (amenah8):

yeah i messaged them hopefully they come XD

OpenStudy (amenah8):

but i'm doing ok with solving so far

OpenStudy (bonnieisflash1.0):

@osprey

OpenStudy (bonnieisflash1.0):

oh you here @osprey

OpenStudy (bonnieisflash1.0):

how did you know she needed help @zepdrix

OpenStudy (amenah8):

ug i got the derivative to be -12(x-18)x, set it equal to 0 and solved for x which I got to be 0

OpenStudy (amenah8):

is that right?

OpenStudy (osprey):

this is another optimisation problem. I'm in danger of getting out my linear programming and other minimising/maximising stuff ... Unf I don't have it to hand (nor to brain). @bonnieisflash1.0 should I say Miaow ? ("Batman returns" is a "nice" film, and "Miss Kitty" steals the show completely.

OpenStudy (bonnieisflash1.0):

go in your nest and hide there @osprey

OpenStudy (osprey):

@bonnieisflash1.0 ouch ! will do

OpenStudy (bonnieisflash1.0):

okay

zepdrix (zepdrix):

-12(x-18)x for derivative? whuuu? Where is 18 coming from??

OpenStudy (amenah8):

I was trying to do what agentsmith said: V=x^2 * y, and length plus girth = 108=4x+y i solved for y using the second equation and got y=108-4x and then plugged that into the Volume equation and took the derivative

OpenStudy (amenah8):

So I was trying to take the derivative of V=X^2 * (108-4x)

zepdrix (zepdrix):

So product rule, yes?

OpenStudy (amenah8):

right

OpenStudy (amenah8):

(x^2)(-4) + (108-4x)(2x)

zepdrix (zepdrix):

ok good

OpenStudy (amenah8):

2(108-4x)x-4x^2 ??

zepdrix (zepdrix):

Put the x in front, it's ugly back there >.< 2x(108-4x)-4x^2 Ok sooooo you've got your derivative of volume. Look for the critical value which will maximize volume.

OpenStudy (amenah8):

wait don't i have to solve for x?

OpenStudy (amenah8):

@zepdrix

zepdrix (zepdrix):

Yes, that's what it means to find critical points, solve for x.